# Thermal evolution of dark matter in the early universe from a symplectic glueball model

### v1

@@ 3 Results
(where $M$ is the mass of the lightest physical state in the scalar channel in continuum, while $\mathcal{C}_{2}(\mbox{fund})$ and $\mathcal{C}_{2}(\mbox{adj})$ respectively denote the eigenvalue of the quadratic Casimir operator for the fundamental and for the adjoint representation of the gauge group, defined in the appendix A), depends only on the number of spacetime dimensions $D$, and is universal for all purely gluonic Yang–Mills theories, taking the value $\eta(4)=5.41(12)$ in the case of four spacetime dimensions. Also, note that for the theory we are considering the $\mathcal{C}_{2}(\mbox{fund})/\mathcal{C}_{2}(\mbox{adj})$ ratio equals $\frac{5}{12}$, which is close to the value it takes for the $\operatorname{\mathrm{SU}}(3)$ Yang–Mills theory, i.e., $\frac{4}{9}$.
As for the behavior of the $\operatorname{\mathrm{Sp}}(2)$ theory at finite temperature, the study reported in ref. [34] already revealed the existence of a first-order deconfining phase transition at a finite critical temperature which, when expressed in units of the square root of the string tension, is $\frac{T_{\mbox{\tiny{c}}}}{\sqrt{\sigma}}=0.6875(18)$.
@@
Our study of the deconfinement temperature of the $\operatorname{\mathrm{Sp}}(2)$ is based on a finite-size scaling of Binder cumulants. The plot in figure 1 shows an example of our results for the behavior of the Binder cumulant associated with the Polyakov loop: the data displayed in the figure were obtained from simulations on a lattice with $N_{\tau}=8$ sites in the Euclidean-time direction. The data are plotted against $\beta=8/g_{0}^{2}$ (where $g_{0}$ is the bare coupling of the theory), and for different values of the extent of the lattice (in units of the lattice spacing) along the spatial directions, which are denoted by symbols of different shapes and colors. The critical value of $\beta$ corresponding to the deconfinement transition can be estimated from the crossing of the curves interpolating the data at different spatial volumes.
@@
$k_{0}=1.892(3),\quad k_{1}=1.371(8),\quad k_{2}=-0.21(4),\quad k_{3}=-0.35(7),$ (3.3)
with $\chi^{2}_{\mbox{\tiny{red}}}\simeq 2.2$; the resulting curve, which is our non-perturbative scale setting for the theory, is shown in fig. 2, alongside the data. We studied how the scale setting varies, if the fitting form on the right-hand side of eq. (3.2) is replaced by a polynomial of different order, and found that the results remain consistent, as the corresponding systematic deviations are less than the statistical uncertainties.
@@
While the results in table 2 allow one to set the value of the lattice spacing for a set of eight values of $\beta$ from the relation $a_{\mbox{\tiny{lat}}}(\beta_{\mbox{\tiny{c}}})=\frac{1}{N_{\tau}T_{\mbox{\tiny{c}}}}$, the interpolation in eq. (3.2) provides a continuous parameterization for the lattice spacing as a function of $\beta$. To extend the correspondence between values of the coupling and values of $a_{\mbox{\tiny{lat}}}$ to finer values of the lattice spacing (which, in particular, is necessary in order to study high temperatures, as well as the continuum limit of the lattice theory), we supplement this parameterization with a two-loop perturbative expression in an improved lattice scheme [166], whereby the lattice spacing can be expressed in terms of a mean-field-improved square coupling $g_{\mbox{\tiny{imp}}}^{2}=\frac{g_{0}^{2}}{\langle P\rangle}$ [167] as follows
@@
To compare the mean-field-improved perturbative scale setting with the non-perturbative one, it is convenient to introduce the parameter $\beta_{\mbox{\tiny{imp}}}=4N/g_{\mbox{\tiny{imp}}}^{2}=\beta\langle P\rangle$. Note that the $\langle P\rangle$ expectation value is a monotonically increasing function of $\beta$, ranging from 00 (at $\beta=0$, i.e., in the strong-coupling limit) to $1$ (for $\beta\to\infty$, namely, in the weak-coupling limit). As a consequence, one always has $\beta_{\mbox{\tiny{imp}}}\leq\beta$; in particular, for the $\beta$ values that we considered in our lattice simulations, $\beta_{\mbox{\tiny{imp}}}$ is in the ballpark of $\frac{\beta}{2}$. Therefore, it is interesting to compare the non-perturbative scale setting with the mean-field-improved one for values of $\beta_{\mbox{\tiny{imp}}}$ of the order of four. An example of such comparison is displayed in the inset plot of fig. 2, where the lattice spacing (in units of the inverse of the critical temperature) is plotted against $\beta_{\mbox{\tiny{imp}}}$: as expected, these two parameterizations of the lattice spacing as a function of the coupling tend to be in better and better agreement when $\beta_{\mbox{\tiny{imp}}}$ is increased, while they deviate from each other at larger values of the coupling.
Having set the scale, it is then possible to study the thermodynamic quantities in a continuous range of temperatures across the deconfinement phase transition, and to carry out their extrapolation to the continuum limit. As an example, figure 3 shows our results for the trace dimensionless ratio of the energy-momentum tensor $\Delta$ over the fourth power of the temperature, as obtained from eq. (2.15): the plot shows the results of our simulations at three different values of the lattice spacing, corresponding to simulations on lattices with $N_{\tau}=5$, $6$, and $7$ sites along the Euclidean-time direction. Note that the trace of the energy-momentum tensor is, in some sense, the most basic equilibrium-thermodynamics quantity that is directly extracted from differences of plaquettes computed in our lattice simulations, while the pressure is obtained by numerical integration of the former, according to eq. (2.13), and the densities of energy and entropy are obtained as linear combinations of $\Delta$ and $p$, through eq. (2.16) and eq. (2.17), respectively.
@@
Figure 3: Trace of the energy-momentum tensor $\Delta$ in the $\operatorname{\mathrm{Sp}}(2)$ theory, expressed in units of the fourth power of the temperature, as obtained from our simulations at finite values of the lattice spacing, equal to one fifth (black squares), one sixth (red triangles), and one seventh (green circles) of the lattice extent in the Euclidean-time direction.
@@
The data shown in figure 3 reveal a behavior consistent with what is observed in other non-Abelian Yang–Mills theories [53, 60, 54, 176, 177], including, in particular, rather mild discretization effects, which are most clearly visible close to the peak of $\frac{\Delta}{T^{4}}$.
@@
It is then possible to derive the results for $\frac{p}{T^{4}}$, for $\frac{\epsilon}{T^{4}}$, and for $\frac{s}{T^{3}}$, and to proceed to their extrapolation to the continuum limit $a_{\mbox{\tiny{lat}}}\to 0$. We carry out the continuum extrapolation of the equilibrium-thermodynamics quantities by first interpolating the values of $\Delta/T^{4}$ at each $N_{\tau}$ using splines (a choice that can be argued to introduce the minimum amount of information into the data), and then extrapolating the values thus obtained at a given temperature as a function of $1/N_{\tau}$. Repeating this procedure at a sufficiently large number of temperatures, one can get an arbitrarily precise approximation of a continuous curve, from which the pressure is then obtained by numerical integration, according to eq. (2.15); finally, the energy density and the entropy density are computed from the extrapolated values of $\Delta/T^{4}$ and $p/T^{4}$ using eq. (2.16) and eq. (2.17). The results of this extrapolation to the continuum limit are shown in figure 4.
@@
Figure 4: Results for thermodynamic-equilibrium quantities extrapolated to the continuum limit, and plotted as a function of the $T/T_{\mbox{\tiny{c}}}$ ratio: the figure shows the entropy density in units of $T^{3}$ (magenta symbols), the energy density (blue symbols), the trace of the energy-momentum tensor (brown symbols), and the pressure (red symbols) in units of $T^{4}$. The horizontal bars on the right-hand side of the figure show, from top to bottom, the values corresponding to the Stefan–Boltzmann limit for $s/T^{3}$, for $\epsilon/T^{4}$, and for $p/T^{4}$.

### v2

@@ 3 Results
(where $M$ is the mass of the lightest physical state in the scalar channel in continuum, while $\mathcal{C}_{2}(\mbox{fund})$ and $\mathcal{C}_{2}(\mbox{adj})$ respectively denote the eigenvalue of the quadratic Casimir operator for the fundamental and for the adjoint representation of the gauge group, defined in the appendix A), depends only on the number of spacetime dimensions $D$, and is universal for all purely gluonic Yang–Mills theories, taking the value $\eta(4)=5.41(12)$ in the case of four spacetime dimensions. Also, note that for the theory we are considering the $\mathcal{C}_{2}(\mbox{fund})/\mathcal{C}_{2}(\mbox{adj})$ ratio equals $5/12$, which is close to the value it takes for the $\operatorname{\mathrm{SU}}(3)$ Yang–Mills theory, i.e., $4/9$.
As for the behavior of the $\operatorname{\mathrm{Sp}}(2)$ theory at finite temperature, the study reported in ref. [35] already revealed the existence of a first-order deconfining phase transition at a finite critical temperature which, when expressed in units of the square root of the string tension, is $T_{\mbox{\tiny{c}}}/\sqrt{\sigma}=0.6875(18)$.
@@
Our study of the deconfinement temperature of the $\operatorname{\mathrm{Sp}}(2)$ theory is based on a finite-size scaling of Binder cumulants. The plot in figure 1 shows an example of our results for the behavior of the Binder cumulant associated with the Polyakov loop: the data displayed in the figure were obtained from simulations on a lattice with $N_{\tau}=8$ sites in the Euclidean-time direction. The data are plotted against $\beta=8/g_{0}^{2}$ (where $g_{0}$ is the bare coupling of the theory), and for different values of the extent of the lattice (in units of the lattice spacing) along the spatial directions, which are denoted by symbols of different shapes and colors. The critical value of $\beta$ corresponding to the deconfinement transition can be estimated from the crossing of the curves interpolating the data at different spatial volumes.
@@
$k_{0}=1.892(3),\quad k_{1}=1.371(8),\quad k_{2}=-0.21(4),\quad k_{3}=-0.35(7).$ (3.3)
This fit, with four degrees of freedom, leads to $\chi^{2}_{\mbox{\tiny{red}}}\simeq 2.2$, corresponding to $p\simeq 0.066$. The resulting curve, which is our non-perturbative scale setting for the theory, is shown in fig. 2, alongside the data. We also studied how the scale setting varies, if the fitting form on the right-hand side of eq. (3.2) is replaced by a polynomial of different order, and found that the results remain consistent, as the corresponding systematic deviations are less than the statistical uncertainties. Note that we chose to express the right-hand side of eq. (3.2) in terms of an expansion around a $\beta_{\star}$ value close to the center of the $\beta$ interval covered by our lattice simulations: this choice is expected to minimize the numerical artifacts affecting polynomial interpolations. [^6] A simple and general argument goes as follows. Consider the approximation of a continuous and differentiable function $f$ on the real finite $x_{\mbox{\tiny{min}}},x_{\mbox{\tiny{max}}}]$ interval through a sequence of polynomials $p_{n}(x-x_{\star})$ of degree $n$. According to Weierstrass’ approximation theorem, the maximum error in the approximation, $\sup_{x\in[x_{\mbox{\tiny{min}}},x_{\mbox{\tiny{max}}}]}|f(x)-p_{n}(x-x_{\star})|$, can be made arbitrarily small (by increasing the polynomial degree $n$), as long as the $x_{\mbox{\tiny{min}}},x_{\mbox{\tiny{max}}}]$ interval lies within the radius of convergence of the expansion about $x_{\star}$. If that is not the case, then the maximum error in the approximation can remain finite—and, in fact, may even become arbitrarily large—regardless of the degree of the polynomial [183]. While approximating an unknown function $f$ (which may have poles in the complex plane at unknown locations), choosing $x_{\star}$ to be the central point of the $x_{\mbox{\tiny{min}}},x_{\mbox{\tiny{max}}}]$ interval minimizes the area of the smallest circle in the complex plane of $x$ in which the real $x_{\mbox{\tiny{min}}},x_{\mbox{\tiny{max}}}]$ interval is entirely contained, therefore maximizing the probability that such circle does not contain any complex poles of $f$ and thus it minimizes the chances that the approximation may have potentially unbounded errors in the considered interval.
@@
While the results in table 2 allow one to set the value of the lattice spacing for a set of eight values of $\beta$ from the relation $a_{\mbox{\tiny{lat}}}(\beta_{\mbox{\tiny{c}}})=1/(N_{\tau}T_{\mbox{\tiny{c}}})$, the interpolation in eq. (3.2) provides a continuous parameterization for the lattice spacing as a function of $\beta$. To extend the correspondence between values of the coupling and values of $a_{\mbox{\tiny{lat}}}$ to finer lattice spacings (which, in particular, is necessary in order to study high temperatures, as well as the continuum limit of the lattice theory), we supplement this parameterization with a two-loop perturbative expression in an improved lattice scheme [184], whereby the lattice spacing can be expressed in terms of a mean-field-improved square coupling $g_{\mbox{\tiny{imp}}}^{2}=g_{0}^{2}/\langle P\rangle$ [185] as follows
@@
To compare the mean-field-improved perturbative scale setting with the non-perturbative one, it is convenient to introduce the parameter $\beta_{\mbox{\tiny{imp}}}=4N/g_{\mbox{\tiny{imp}}}^{2}=\beta\langle P\rangle$. Note that the $\langle P\rangle$ expectation value is a monotonically increasing function of $\beta$, ranging from 0 (at $\beta=0$, i.e., in the strong-coupling limit) to $1$ (for $\beta\to\infty$, namely, in the weak-coupling limit). As a consequence, one always has $\beta_{\mbox{\tiny{imp}}}\leq\beta$; in particular, for the $\beta$ values that we considered in our lattice simulations, $\beta_{\mbox{\tiny{imp}}}$ is in the ballpark of $\beta/2$. Therefore, it is interesting to compare the non-perturbative scale setting with the mean-field-improved one for values of $\beta_{\mbox{\tiny{imp}}}$ of the order of four. An example of such comparison is displayed in the inset plot of fig. 2, where the lattice spacing (in units of the inverse of the critical temperature) is plotted against $\beta_{\mbox{\tiny{imp}}}$: as expected, these two parameterizations of the lattice spacing as a function of the coupling tend to be in better and better agreement when $\beta_{\mbox{\tiny{imp}}}$ is increased, while they deviate from each other at larger values of the coupling. It is interesting to discuss the impact of the systematic uncertainty induced by combining a scale setting obtained from the interpolation of simulation results (at intermediate $\beta$ values), and the fixed-order perturbative expression in eq. (3.4) (in the $\beta>8.025$ region). In particular, one may suspect that the matching between a genuinely non-perturbative (and thus, in a sense, “all-order”) scale setting and a second-order perturbative formula, imposing that they take the same value at the point where they are matched, could be unjustified, as it corresponds to neglecting the impact of higher-order perturbative contributions. It should be noted, however, that eq. (3.4) is an expansion in terms of the improved lattice coupling $g_{\mbox{\tiny{imp}}}^{2}=g_{0}^{2}/\langle P\rangle$, which can be interpreted as the “effective coupling” experienced by a background field in the lattice theory (in a mean-field approximation) [185]: since the definition of $g_{\mbox{\tiny{imp}}}^{2}$ involves the plaquette expectation value, which we determined non-perturbatively from lattice simulations, the expression in eq. (3.4) is, in fact, expected to encode some non-perturbative information and to satisfy the key properties expected for a “good” coupling for a lattice gauge theory—including, in particular, satisfactory convergence properties in perturbation theory [194]. For $\operatorname{\mathrm{SU}}(N)$ gauge theories, it was shown in ref. [184] that a fixed-order scale setting in terms of $g_{\mbox{\tiny{imp}}}$ is in remarkably good agreement with a fully non-perturbative scale setting defined in terms of the Schrödinger functional [195, 196] over a surprisingly wide range of couplings, so that the matching between the two different parametrizations for the lattice scale as a function of $\beta$ is actually smooth. In principle, this should be tested for the $\operatorname{\mathrm{Sp}}(2)$ gauge theory, too, by means of a dedicated lattice calculation. We expect that the computational costs for such study could be comparable with those of analogous calculations that were carried out for the $\operatorname{\mathrm{SU}}(3)$ theory (for a recent, high-precision study, see ref. [197]) and for the $\operatorname{\mathrm{SU}}(4)$ theory [198]: the dimension of the Lie algebra of the $\operatorname{\mathrm{Sp}}(2)$ theory ($10$) is intermediate between those of these two $\operatorname{\mathrm{SU}}(N)$ theories ($8$ and $15$), the size of the gauge-group matrices for the $\operatorname{\mathrm{Sp}}(2)$ lattice gauge theory ($4$) is the same as for the $\operatorname{\mathrm{SU}}(4)$ theory, while the rank of the Lie algebra of the symplectic theory ($2$) is the same as that of $\operatorname{\mathrm{SU}}(3)$ (even though the “fundamental domain” involved in the formulation of the Schrödinger-functional is, of course, different, since it depends on the weights of the Lie algebra). For our present purposes, however, we limit ourselves to a very crude, order-of-magnitude estimate of the systematic uncertainty on $\Lambda_{\mbox{\tiny{imp}}}$ from the matching procedure. In particular, we note that, at the matching point (corresponding to $\beta=8.025$, or to $\beta_{\mbox{\tiny{imp}}}\simeq 5.075$) the difference in slope of the tangent to our fully non-perturbative scale setting and to the improved two-loop expression in eq. (3.4) is tiny, $0.017$, suggesting that the effect of residual higher-order corrections (which are expected to contribute to the slope of the curve) is small; at $\beta$ values larger than $8.025$, such corrections are expected to be even smaller (as they originate from higher powers of the coupling). The results for the equation of state that we present in this section are limited to temperatures relatively close to $T_{\mbox{\tiny{c}}}$: in particular, we report continuum results for the pressure, energy density, and entropy density at temperatures up to $T=2T_{\mbox{\tiny{c}}}$, extrapolating from simulation results up to $N_{\tau}=7$. This means that the finest lattice spacing that we considered in our calculations of the equation of state is $1/(2\cdot 7T_{\mbox{\tiny{c}}})$, that is about $70\%$ of the finest lattice spacing that we determined non-perturbatively (which is the one corresponding to the critical temperature for $N_{\tau}=10$, namely $1/(10T_{\mbox{\tiny{c}}})$), and corresponds to $\beta_{\mbox{\tiny{imp}}}\simeq 5.48$, i.e., is only slightly out of the range of our non-perturbative scale setting. Assuming (as a conservative upper bound) that the deviation between the improved perturbative scale setting and a fully non-perturbative scale setting can be described as a linear function with slope $0.017$ (and considering that at $\beta=8.025$ we have $a_{\mbox{\tiny{lat}}}T_{\mbox{\tiny{c}}}=1/10$) we arrive at the crude estimate that the largest relative uncertainty induced by our scale-setting procedure is of the order of $7\%$.
As a final comment on the scale-setting procedure, one can also note that the $\operatorname{\mathrm{Sp}}(2)$ lattice gauge theory at zero temperature has no bulk phase transition [35], and this is generally expected to make the scale-setting procedure and the continuum limit both conceptually and practically more straightforward than, for example, in $\operatorname{\mathrm{SU}}(N)$ gauge theories with more than three color charges [199]. In any case, the range of couplings that we investigated in the present work is far from the strong-coupling regime (corresponding to $\beta\to 0$), in which lattice artifacts are expected to be significant.
Having set the scale, it is then possible to study the thermodynamic quantities in a continuous range of temperatures across the deconfinement phase transition, and to carry out their extrapolation to the continuum limit. As an example, figure 3 shows our results for the dimensionless ratio of the trace of the energy-momentum tensor $\Delta$ to the fourth power of the temperature, as obtained from eq. (2.15): the plot shows the results of our main set of simulations at three different values of the lattice spacing, corresponding to simulations on lattices with $N_{\tau}=5$, $6$, and $7$ sites along the Euclidean-time direction. In addition, in order to check that the results at these values of the lattice spacing are already reasonably close to the continuum limit, the plot also includes results obtained from simulations at five temperatures close to the deconfinement transition ($T/T_{\mbox{\tiny{c}}}=0.9$, $0.95$, 11, $1.05$, and $1.1$) from simulations on lattices with $N_{\tau}=8$ and $9$ sites along the Euclidean-time direction, which, for the sake of clarity, are denoted by larger symbols in the figure. The results obtained from these finer lattices appear to be consistent with those obtained from the three largest ensembles (namely, those with $N_{\tau}=5$, $6$, and $7$). It is worth noting that the computational costs of the simulations grow quite quickly, when $N_{\tau}$ is increased. The reason for this can be appreciated by looking at eq. (2.15): the trace of the energy-momentum tensor is proportional to the difference of the plaquette expectation values at finite $T$ and at $T=0$ and to the fourth power of $N_{\tau}$. This means that the $(\langle P\rangle_{T}-\langle P\rangle_{0})$ difference scales (approximately) as $N_{\tau}^{-4}$, and if the signal-to-noise ratio in the determination of $\Delta/T^{4}$ has to be held fixed, then the uncertainties on $\langle P\rangle_{T}$ and $\langle P\rangle_{0}$ should also be scaled as $N_{\tau}^{-4}$. In turn, this would require a number of independent lattice configurations growing as $N_{\tau}^{8}$. These are not the only technical challenges that should be faced when $N_{\tau}$ grows. It is known that lattice simulations on very fine lattices incur severe ergodicity problems, due to the phenomenon of critical slowing down [200]; these include, in particular, the inability to efficiently probe sectors of different topological charge [201]. Even though these technical reasons limited the precision that we could obtain in our simulations of the $\operatorname{\mathrm{Sp}}(2)$ theory on the finest lattices, we note that the general consistency of the lattice results for the $N_{\tau}$ values that we considered in this work has also been observed elsewhere for $\operatorname{\mathrm{SU}}(N)$ gauge theories [202, 64]. This gives confidence on the validity of our extrapolations to the continuum limit, within the uncertainties reported here (which are larger than those of previous $\operatorname{\mathrm{SU}}(3)$ calculations). Note that the trace of the energy-momentum tensor is, in some sense, the most basic equilibrium-thermodynamics quantity that is directly extracted from differences of plaquettes computed in our lattice simulations, while the pressure is obtained by numerical integration of the former, according to eq. (2.13), and the densities of energy and entropy are obtained as linear combinations of $\Delta$ and $p$, through eq. (2.16) and eq. (2.17), respectively.
@@
Figure 3: Trace of the energy-momentum tensor $\Delta$ in the $\operatorname{\mathrm{Sp}}(2)$ theory, expressed in units of the fourth power of the temperature, as obtained from our simulations at finite values of the lattice spacing, equal to one fifth (black squares), one sixth (red triangles pointing up), and one seventh (green circles) of the lattice extent in the Euclidean-time direction. In addition, the figure also includes the results obtained from simulations on lattices with values of the lattice spacing equal to one eighth (blue triangles pointing down) and one ninth (orange diamonds) of the lattice extent in the Euclidean-time direction, for a more limited set of temperatures ($T/T_{\mbox{\tiny{c}}}=0.9$, $0.95$, 11, $1.05$, and $1.1$).
@@
The data shown in figure 3 reveal a behavior consistent with what is observed in other non-Abelian Yang–Mills theories [63, 70, 64, 203, 204], including, in particular, rather mild discretization effects, which are most clearly visible close to the peak of $\Delta/T^{4}$.
@@
It is then possible to derive the results for $p/T^{4}$, for $\epsilon/T^{4}$, and for $s/T^{3}$, and to proceed to their extrapolation to the continuum limit $a_{\mbox{\tiny{lat}}}\to 0$. We carry out the continuum extrapolation of the equilibrium-thermodynamics quantities by first interpolating the values of $\Delta/T^{4}$ at each $N_{\tau}$ using splines (a choice that is expected to yield optimal bounds on the interpolation uncertainties [205]), and then extrapolating the values thus obtained at a given temperature as a function of $1/N_{\tau}^{2}$. Repeating this procedure at a sufficiently large number of temperatures, one can get an arbitrarily precise approximation of a continuous curve, from which the pressure is then obtained by numerical integration, according to eq. (2.15); finally, the energy density and the entropy density are computed from the extrapolated values of $\Delta/T^{4}$ and $p/T^{4}$ using eq. (2.16) and eq. (2.17).
While the precision of our continuum-extrapolated results is worse than, for example, the precision that has been achieved in the past few years for the equation of state in the pure-glue sector of QCD, it may be instructive to consider the quality of the extrapolation at some reference temperatures. As an example, in figure 4 we show the extrapolation to the continuum limit of the large-volume extrapolation of our results for $\Delta/T^{4}$ at $T/T_{\mbox{\tiny{c}}}=1.5$ (denoted by the black circles), at the three values of the lattice spacing for which we could obtain sufficiently precise simulation results, corresponding to $N_{\tau}=5$, $N_{\tau}=6$, and $N_{\tau}=7$. As the leading finite-cutoff corrections in the lattice discretization with Wilson’s action are proportional to $a_{\mbox{\tiny{lat}}}^{2}$, the data are fitted to a constant plus a term linear in $1/N_{\tau}^{2}$. The fit (shown by the dashed black line) yields $1.867(12)$ for the intercept (denoted by the red square), while the result for the slope, 0.2(4)-0.2(4), is compatible with zero. The reduced $\chi^{2}$ of the fit is $0.21$, with one degree of freedom, and the non-diagonal elements of the correlation matrix of the fit parameters are 0.947-0.947.
Figure 4: Continuum extrapolation for $\Delta/T^{4}$ at $T/T_{\mbox{\tiny{c}}}=1.5$: the data obtained from the infinite-volume extrapolation of the results of simulations at $T/T_{\mbox{\tiny{c}}}=1.5$ on lattices with $N_{\tau}=5$, $N_{\tau}=6$, and $N_{\tau}=7$ points along the Euclidean-time direction (black circles) are fitted as a function of $1/N_{\tau}^{2}$, including a constant and linear term. The fit result is denoted by the dashed black line. The red square denotes the extrapolation to the continuum limit.
Similar fit results are obtained at the other temperatures that we studied in this work. Considering, for example, some selected temperatures in different parts of the $T/T_{\mbox{\tiny{c}}}$ range that we considered, namely $T/T_{\mbox{\tiny{c}}}=0.95$ (in the confining phase), $T/T_{\mbox{\tiny{c}}}=1.15$ (in the region of the peak of $\Delta/T^{4}$), and $T/T_{\mbox{\tiny{c}}}=2$ (the highest temperature at which we carried out the continuum extrapolation), we respectively obtain $\chi^{2}_{\mbox{\tiny{red}}}=0.19$, $\chi^{2}_{\mbox{\tiny{red}}}=0.18$, and $\chi^{2}_{\mbox{\tiny{red}}}=3.81$. The slope of the interpolating straight line as a function of $1/N_{\tau}^{2}$ is compatible with zero, with the exception of the $T/T_{\mbox{\tiny{c}}}=1.15$ case (for which it is found to be $14.5(6)$).
The results of the extrapolation to the continuum limit are shown in figure 5, while in tables 3 and 4 we report a breakdown of the error budget for $\Delta/T^{4}$ and for $p/T^{4}$, respectively, at the selected temperatures we mentioned above (while we omit the corresponding discussion for $\epsilon/T^{4}$ and for $s/T^{3}$, as the latter quantities were obtained from linear combinations of the former), which deserves some comments. First of all, the various contributions to the total uncertainties are somewhat entangled with each other: for example, given the results for $\Delta/T^{4}$ obtained at a given $N_{\tau}$ from three different lattice volumes, with their own (typically: similar) uncertainties, it is somewhat ambiguous to state what is the “statistical contribution” to the final error budget; for this reasons, the quoted figures should be considered as order-of-magnitude estimates of the various contributions to the total error budget, whose main purpose is to highlight which are the bottlenecks affecting the precision of the final extrapolated results. A second observation is that usually the extrapolation to the thermodynamic limit is quite smooth: the simulations carried out for this work were performed on lattices with space-time cross-sections characterized by aspect ratios $LT\simeq 4$, which typically is enough to guarantee a sufficiently strong suppression of finite-volume corrections. This is a consequence of the fact that, as table 1 shows, in the confining phase the purely gluonic $\operatorname{\mathrm{Sp}}(2)$ theory has a large mass gap: finite-volume corrections are exponentially suppressed with a characteristic length scale given by the Compton wavelength associated with the lightest physical state, which, in natural units, is approximately five times smaller than the inverse of the critical temperature, and thus approximately $20$ times smaller than the linear extent of the lattice in each of the three main spatial directions. In the deconfined phase, on the other hand, finite-volume corrections are suppressed by the fact that the deconfined plasma has a finite screening length. For this reason, some of the $\Delta/T^{4}$ values computed numerically at a fixed $N_{\tau}$ and from lattices of different volumes turn out to be fully compatible with each other within their uncertainties: in this case, carrying out a thermodynamic extrapolation by fitting those values to a constant plus a term inversely proportional to the volume reveals that the interpolating straight line is flat, and, in practice, the extrapolated constant is numerically close to the weighted average of the $\Delta/T^{4}$ values obtained from the simulations at the different volumes. Thirdly, the uncertainties associated with the scale setting—which, beside $T/T_{\mbox{\tiny{c}}}$, directly affect the evaluation of $\Delta/T^{4}$, because the latter is computed according to eq. (2.15), which involves the relation between the lattice spacing and $\beta$—are typically very small: this is mostly due to the good precision of our determination of the non-perturbative scale setting according to eq. (3.2), where especially the leading terms $k_{0}$ and $k_{1}$ in eqs. (3.3) have very small uncertainties. Next, we also note that the uncertainties associated with the spline interpolations are always very small, too: in this case, this is due to the fact that in our simulations we sampled a very dense set of temperatures in the range of interest, and (except in the proximity of $T_{\mbox{\tiny{c}}}$) the behavior of $\Delta/T^{4}$ as a function of $T/T_{\mbox{\tiny{c}}}$ is quite smooth. The fineness of the grid of temperature values that was used in the numerical integration of $p/T^{4}$ implies that also the uncertainties associated with this quantity are rather small. In practice, it is only close to $T_{\mbox{\tiny{c}}}$ (i.e., in the narrow temperature range where the variation of $\Delta/T^{4}$ is fastest) that the systematics associated with the numerical integration are not completely negligible. For $p/T^{4}$, it should be noted that the uncertainties tend to accumulate (more precisely: to sum up in quadrature) in the integration process; however, the highly improved numerical scheme that we used, which is defined in eq. (2), allowed us to keep them under control. Finally, the extrapolation to the continuum limit (carried out fitting the results extrapolated to the infinite-volume limit for each $N_{\tau}$ to a constant plus a term proportional to $N_{\tau}^{-2}$) was found to have a non-negligible impact on the total uncertainty of $\Delta/T^{4}$, in particular for temperatures corresponding to the maximum of this quantity. Attempting to include also an $N_{\tau}^{-4}$ correction (at least at the very few temperatures for which we also had results for $\Delta/T^{4}$ from simulations with $N_{\tau}=8$ and $N_{\tau}=9$) did not lead to any improvement in the results, nor on the estimate of the continuum-extrapolation uncertainties.
@@
Figure 5: Results for thermodynamic-equilibrium quantities extrapolated to the continuum limit, and plotted as a function of the $T/T_{\mbox{\tiny{c}}}$ ratio: the main plot shows the entropy density in units of $T^{3}$ (magenta symbols), the energy density (blue symbols), the trace of the energy-momentum tensor (brown symbols), and the pressure (red symbols) in units of $T^{4}$. The horizontal bars on the right-hand side of the figure show, from top to bottom, the values corresponding to the Stefan–Boltzmann limit for $s/T^{3}$, for $\epsilon/T^{4}$, and for $p/T^{4}$. In the inset plot, the same quantities (except for $\Delta/T^{4}$, whose value in the free limit is zero) are normalized by their values in the Stefan–Boltzmann limit, with the same colors as in the main plot.
——- $T/T_{\mbox{\tiny{c}}}$ notes $\Delta/T^{4}$ uncertainty on $\Delta/T^{4}$ $0.95$ confining phase $0.2115$ $0.0026~\left\{\begin{array}[]{l}0.002~\mbox{statistics}\\ 0.0005~\mbox{$V\to\infty$ extrap.}\\ 0.0002~\mbox{scale-setting}\\ <0.0001~\mbox{spline interp.}\\ 0.001~\mbox{$a_{\mbox{\tiny{lat}}}\to 0$ extrap.}\end{array}\right$. $1.15$ peak of $\Delta/T^{4}$ $2.715$ $0.015~\left\{\begin{array}[]{l}0.013~\mbox{statistics}\\ 0.006~\mbox{$V\to\infty$ extrap.}\\ 0.0002~\mbox{scale-setting}\\ 0.0001~\mbox{spline interp.}\\ 0.006~\mbox{$a_{\mbox{\tiny{lat}}}\to 0$ extrap.}\end{array}\right$. $1.5$ $1.867$ $0.012~\left\{\begin{array}[]{l}0.01~\mbox{statistics}\\ 0.003~\mbox{$V\to\infty$ extrap.}\\ 0.0002~\mbox{scale-setting}\\ <0.0001~\mbox{spline interp.}\\ 0.001~\mbox{$a_{\mbox{\tiny{lat}}}\to 0$ extrap.}\end{array}\right$. 22 largest $T$ of our cont. extrap. $1.094$ $0.037~\left\{\begin{array}[]{l}0.028~\mbox{statistics}\\ 0.01~\mbox{$V\to\infty$ extrap.}\\ 0.005~\mbox{scale-setting}\\ <0.0001~\mbox{spline interp.}\\ 0.006~\mbox{$a_{\mbox{\tiny{lat}}}\to 0$ extrap.}\end{array}\right$. ——-
Table 3: Breakdown of the uncertainty budget on the continuum-extrapolated values of $\Delta/T^{4}$ at a selection of temperatures.
[…]
Table 4: Same as in table 3, but for $p/T^{4}$, which, as discussed in the text, is obtained by numerical integration of the continuum-extrapolated results for $\Delta/T^{4}$. Note that the uncertainties on $\Delta/T^{4}$ have a cumulative effect on $p/T^{4}$.